OpenStudy (anonymous):

Use a cross products to solve the proportion. 1. 16/48 = v/12

OpenStudy (raden):

first it would be easier, if we simplify of 16/48 be 1/3. now we have 1/3 = v/12 cross multiplify, get 3v = 12 solve for v ....

OpenStudy (anonymous):

ANSWER IS V=4?

OpenStudy (raden):

yes, you are right